# Multi-Grained Query-Guided Set Prediction Network for Grounded Multimodal Named Entity Recognition

### v1

@@ 3.5. Query-guided Fusion Net
Similarity-aware Aggregator. To mitigate the noise caused by misaligned candidate regions and entity spans, we propose a similarity-aware aggregator to learn fine-grained token-wise alignment between query tokens and regional features/textual tokens. We obtain the output vectors from the query-text cross-attention interaction module and query-region prefix integration module: denoted as $\tilde{H}_{Q}$, $\tilde{H}_{T}$, and $\tilde{H}_{V}$, respectively corresponding to the query set feature, textual feature, and visual feature. We compute the token-wise similarity matrix of the $i$-th query token as follows:
(8) $\alpha_{j}^{\phi}=\frac{\exp{(\tilde{H}_{\phi_{j}}\cdot\tilde{H}_{Q}})}{\sum_{j}\exp{(\tilde{H}_{\phi_{j}}\cdot\tilde{H}_{Q}})}$
@@
where $\phi\in{\{V,T\}}$ represents the visual or textual feature. $\tilde{H}_{\phi_{j}}$ refers to the representation of the $j$-th visual regions or textual tokens. Finally, a fine-grained fusion module is proposed to integrate similarity-aware visual or textual hidden states into the query hidden states:
(9) $\mathrm{F}(\tilde{H}_{Q})=\mathrm{Tanh}(\tilde{H}_{Q}W_{1}+\sum_{\phi}\sum_{j}\lambda_{\phi}\alpha_{j}^{\phi}\tilde{H}_{\phi_{j}})W_{2}+\mathbf{b}$
@@
3.6. Multimodal Set Prediction Network
Instead of relying on the predefined decoding order in autoregressive models, we propose Multimodal Set Prediction Network to maintain suitable inter-entity relationships in a global matching perspective. Specifically, we first feed $u$ query set features associated with visual and textual features into three prediction heads to obtain $u$ predicted quadruple set. Subsequently, we apply Hungarian Algorithm (Kuhn, 1955) to figure out one-to-one matching between the predicted quadruple set and the gold set. The main difficulty of set prediction in training is to evaluate the predicted quadruples according to the gold quadruples. To solve this problem, we design a loss function based on optimal bipartite matching. The details are discussed below.
Span Boundary Localization. Given the output textual representation $\hat{H}_{T}\in\mathbb{R}^{n\times h}$ and the query set representation $\hat{H}_{Q}\in\mathbb{R}^{u\times h}$ from section 3.5, we first expand their dimensions as $\hat{H}_{T}\in\mathbb{R}^{1\times n\times h}$ and $\hat{H}_{Q}\in\mathbb{R}^{u\times 1\times h}$, and then integrate them into a joint representation:
(10) $\mathrm{S^{b}}=\mathrm{ReLU}(\hat{H}_{Q}W_{Q}^{b}+\hat{H}_{T}W_{T}^{b})$
@@ 3.6. Multimodal Set Prediction Network
(11) $P^{s}=\mathrm{sigmoid}(\mathrm{S}_{b}W_{s}^{b})$
@@
(12) $\mathrm{S^{r}}=\mathrm{ReLU}(\hat{H}_{Q}W_{Q}^{r}+\hat{H}_{V}W_{V}^{r}),$
@@
(13) $\mathrm{S^{c}}=\mathrm{ReLU}\left(\left[\hat{H}_{Q}W_{Q}^{c};P^{s}\hat{H_{T}};P^{e}\hat{H_{T}};P^{r}\hat{H}_{V}\right]\right)$
(14) $P^{c}=\mathrm{sigmoid}(\mathrm{S^{c}}W_{s}^{c})$
@@
Bipartite Matching Loss. In this approach, we design a loss function based on optimal bipartite matching. However, since the number of queries $u$ is greater than the total quantity of gold samples $m$, we define a label $\varnothing$ for those unassignable predictions and then add the label $\varnothing$ to the gold set $Y$, replicating $Y$ until it reaches the size of $u$. Our training objective is to find the optimal assignment $\pi$ that minimizes the global cost. Searching for a permutation of $u$ elements $\pi\in\mathbb{O}_{u}$ can be formalized as follows:
(15) $\hat{\pi}=\arg\min_{\pi\in\mathbb{O}_{u}}\sum_{i}^{u}\mathcal{L}_{\mathrm{cost}}\left(Y_{i},\hat{Y}_{\pi(i)}\right),$
@@
where $\mathcal{L}_{\mathrm{cost}}(\cdot)$ denotes the pair matching cost between the gold quadruple $Y_{i}$ and a prediction with index $\pi(i)$. ${1}_{\{\delta\}}$ represents the indicator function that takes $1$ when $\delta$ is true and 0 otherwise. The optimal assignment $\pi$ can then be efficiently solved by the off-the-shelf Hungarian Algorithm (Kuhn, 1955). To jointly train the model, the final loss function is defined as follows:
(16) $\mathcal{L}(Y,\hat{Y})$ $=\sum_{i=1}^{u}\Big{\{}-\log p_{\hat{\pi}(i)}^{c}(Y_{i}^{t})$
@@
Table 1. The statistics of two GMNER Twitter datasets.

### v2

@@ Introduction
Similarity-aware Aggregator. To mitigate the noise caused by misaligned candidate regions and entity spans, we propose a similarity-aware aggregator to learn fine-grained token-wise alignment between query tokens and regional features/textual tokens. We denote the query set, textual, and visual features as $\tilde{H}_{Q}$, $\tilde{H}_{T}$, and $\tilde{H}_{V}$, respectively. We compute the token-wise similarity matrix of the $i$-th query token as follows:
$\alpha_{j}^{\phi}=\frac{\exp{(\tilde{H}_{\phi_{j}}\cdot\tilde{H}_{Q}})}{\sum_{j}\exp{(\tilde{H}_{\phi_{j}}\cdot\tilde{H}_{Q}})}$ (4)
@@
where $\phi\in{\{V,T\}}$ represents the visual or textual feature. $\tilde{H}_{\phi_{j}}$ refers to the representation of the $j$-th visual regions or textual tokens. Finally, the fine-grained fusion for integrating similarity-aware visual or textual hidden states into the query hidden states can be represented as:
$\mathrm{F}(\tilde{H}_{Q})=\mathrm{Tanh}(\tilde{H}_{Q}W_{1}+\sum_{\phi}\sum_{j}\lambda_{\phi}\alpha_{j}^{\phi}\tilde{H}_{\phi_{j}})W_{2}+\mathbf{b}$ (5)
@@
Multimodal Set Prediction Network
Instead of relying on the predefined decoding order in previous methods, we propose a Multimodal Set Prediction Network to maintain suitable inter-entity relationships in a global matching perspective.
Span Boundary Localization. Given the output textual representation $\hat{H}_{T}\in\mathbb{R}^{n\times h}$ and the query set representation $\hat{H}_{Q}\in\mathbb{R}^{u\times h}$ from QFNet, we first expand their dimensions as $\hat{H}_{T}\in\mathbb{R}^{1\times n\times h}$ and $\hat{H}_{Q}\in\mathbb{R}^{u\times 1\times h}$, and then integrate them into a joint representation:
$\mathrm{S^{b}}=\mathrm{ReLU}(\hat{H}_{Q}W_{Q}^{b}+\hat{H}_{T}W_{T}^{b})$ (6)
@@
$P^{s}=\mathrm{sigmoid}(\mathrm{S}_{b}W_{s}^{b})$ (7)
@@
$\mathrm{S^{r}}=\mathrm{ReLU}(\hat{H}_{Q}W_{Q}^{r}+\hat{H}_{V}W_{V}^{r}),$ (8)
@@
Table 1: Performance comparison of different competitive approaches on Twitter-GMNER datasets. Bold represents the optimal result, and underlined represents the suboptimal result. For the baseline methods, $\clubsuit$ indicates that the results are reproduced according to the corresponding papers, and others are from (Yu et al. 2023).
[…]
@@
$\mathrm{S^{c}}=\mathrm{ReLU}\left(\left[\hat{H}_{Q}W_{Q}^{c};P^{s}\hat{H_{T}};P^{e}\hat{H_{T}};P^{r}\hat{H}_{V}\right]\right)$ (9)
$P^{c}=\mathrm{sigmoid}(\mathrm{S^{c}}W_{s}^{c})$ (10)
@@
Bipartite Matching Loss. Following previous works (Shen et al. 2022; Tan et al. 2021; Chen et al. 2024), we construct a loss function based on optimal bipartite matching. However, since the number of queries $u$ is greater than the total quantity of gold samples $m$, we define a label $\varnothing$ for those unassignable predictions and then add the label $\varnothing$ to the gold set $Y$, replicating $Y$ until it reaches the size of $u$. Our training objective is to find the optimal assignment $\pi$ that minimizes the global cost. Searching for a permutation of $u$ elements $\pi\in\mathbb{O}_{u}$ can be formalized as follows:
$\hat{\pi}=\arg\min_{\pi\in\mathbb{O}_{u}}\sum_{i}^{u}\mathcal{L}_{\mathrm{cost}}\left(Y_{i},\hat{Y}_{\pi(i)}\right),$ (11)
@@
where $\mathcal{L}_{\mathrm{cost}}(\cdot)$ denotes the pair matching cost between the gold quadruple $Y_{i}$ and a prediction with index $\pi(i)$. ${1}_{\{\delta\}}$ represents the indicator function that takes $1$ when $\delta$ is true and 0 otherwise. The optimal assignment $\pi$ can then be efficiently solved by the off-the-shelf Hungarian Algorithm (Kuhn 1955). To jointly train the model, the final loss function is defined as follows:
$\mathcal{L}(Y,\hat{Y})$ $=\sum_{i=1}^{u}\Big{\{}-\log p_{\hat{\pi}(i)}^{c}(Y_{i}^{t})$ (12)
@@
Experiments
Experiment Settings
GMNER Datasets. We conduct experiments on the Twitter-GMNER (Yu et al. 2023) and Twitter-FMNERG (Wang et al. 2023). Please refer to Appendix for detailed information about these datasets.
Evaluation Metrics. Following previous works (Yu et al. 2023), we use Precision (Pre.), Recall (Rec.), and F1 Score (F1) as the evaluation metric for GMNER and its two subtasks. See Appendix for a more detailed definition.
Implementation Details All experiments are implemented on 4 NVIDIA RTX3090 GPUs with Pytorch 1.9.1. For a fair comparison, we use the pre-trained BERT-based model [^2] https://huggingface.co/google-bert/bert-base-uncased as the textual encoder, ViT-B/32 from pre-training CLIP [^3] https://huggingface.co/openai/clip-vit-base-patch32 as the visual encoder, and VinVL [^4] https://github.com/pzzhang/VinVL as a class-agnostic RPN. For model’s efficiency, we freeze the visual encoder and RPN and assign the layer of the QFNet as $L=3$. We initialize the learnable query part with a normal distribution $\mathcal{N}(0,0.02)$. For training, we set the batch size to 16, the learning rate to $2\times 10^{-5}$, and the training epoch to 50. Our model uses an Adam optimizer with a linear warmup of ratio 0.05. We tune the optimal query number $u$ to 60 and 120 for Twitter-GMNER and Twitter-FMNERG datasets, respectively. To allow the multi-grained queries to learn query semantics initially, we first train the model for 5 epochs with the freezing parameters of the pre-training model. Please refer to Appendix for baseline methods.
Table 2: F1 scores of previous SOTA methods H-Index, TIGER, and our MQSPN on Twitter-FMNERG datasets for each fine-grained entity type in GMNER task.
Entity Type H-Index TIGER MQSPN Person 45.13 43.78 48.64 Location 62.33 67.69 71.03 Building 32.88 40.00 44.78 Organization 46.68 46.75 49.03 Product 28.19 27.38 29.46 Art 38.89 43.27 42.84 Event 45.56 48.39 51.88 Other 41.81 48.28 43.49 All 46.55 47.20 48.57
Overall Performance
Performance on GMNER, MNER, and EEG. Following (Yu et al. 2023), we also report two subtasks of GMNER, i.e., Multimodal Named Entity Recognition (MNER) and Entity Extraction & Grounding (EEG). MNER aims to identify entity span and entity type pairs, while EEG aims to extract entity span and entity region pairs. Table 1 shows the performance comparison of our method with competitive baseline models on Twitter-GMNER benchmarks.
First, unified models are significantly superior to pipeline methods due to the joint training of MNER and EEG to mitigate error propagation. Second, MQSPN significantly outperforms the MRC-based method MNER-QG by $+4.85\%$ F1 scores. Furthermore, compared with the previous state-of-the-art (SOTA) generative model TIGER, our method MQSPN exhibits superior performance, achieving $+0.15\%$, $+1.14\%$, and $+2.13\%$ F1 scores improvements on the MNER, EEG, and GMNER tasks, respectively.
We attribute the performance improvements of MQSPN to the following reasons: (1) Compared with the sequence generation-based approach H-Index and TIGER, MQSPN eliminates the dependencies on predefined decoding order and establishes suitable inter-entity relationships from a global matching view. (2) Compared with the MRC-based approach MNER-QG, MQSPN can learn fine-grained entity semantics and model distinct intra-entity connections between regions and entities.
Performance on fine-grained GMNER. We validate the fine-grained GMNER ability of different methods on Twitter-FMNERG datasets, and the experimental results are shown in Table 2. It reveals that our proposed MQSPN achieves the best results on all fine-grained entity types except Other and Art among the state-of-the-art models. Specifically, we achieve a great improvement of $+3.51\%$, $+3.34\%$, $+4.78\%$, $+2.28\%$, $+1.27\%$, $+3.49\%$, and $+1.37\%$ F1 scores on the Person, Location, Building, Organization, Product, Event, and All entity types respectively. The experimental results demonstrate that through robust modeling of intra-entity and inter-entity relationships, MQSPN exhibits superior capability in fine-grained multimodal entity understanding.
Table 3: Ablation study of each component on overall F1 score of different tasks.
Module Settings Twitter-GMNER Twitter-FMNERG GMNER MNER EEG GMNER MNER EEG MQSPN Default 58.76 80.43 62.40 48.57 67.09 62.64 MQS w/o PTQ 56.95 ($\downarrow$1.81) 79.78 61.51 47.23 ($\downarrow$1.34) 65.86 61.75 w/o LEQ 56.02 ($\downarrow$2.74) 78.64 60.30 45.39($\downarrow$3.18) 62.91 59.39 QFNet w/o QCT 57.21 ($\downarrow$1.55) 80.04 61.88 47.61($\downarrow$0.96) 66.14 62.24 w/o QPI 57.86 ($\downarrow$0.90) 80.11 61.91 48.33($\downarrow$0.24) 66.59 62.55 w/o SAG 58.13 ($\downarrow$0.63) 80.23 62.26 47.96($\downarrow$0.61) 66.37 62.37 MSP w/o BML 56.58 ($\downarrow$2.18) 79.26 61.37 46.81($\downarrow$1.76) 64.43 60.81
Ablation Study
Ablation setting. To verify the effectiveness of each designed component in MQSPN, we conduct a comprehensive ablation study on the Twitter-GMNER and Twitter-FMNERG dataset: For Multi-grained Query Set (MQS), (1) w/o PTQ: we remove the prompt-based type-grained part of the query set. (2) w/o LEQ: we replace the learnable queries part with human-designed queries (the same as MNER-QG query construction). For Query-guided Fusion Net (QFNet), (3) w/o QCT: we encode sentences and the query set using the original BERT without query-text cross-attention. (4) w/o QPI: we eliminate the query-region prefix integration. (5) w/o SAG: we eliminate the similarity-aware aggregator. For Multimodal Set Prediction (MSP), (6) w/o BML: we replace the bipartite matching loss with a joint cross-entropy loss in the fixed permutation of the entities. The experimental results are shown in Table 3.
The effectiveness of Multi-grained Query Set. In Table 3, we observe a clear F1 scores drop in model performance (1.81% in Twitter-GMNER and 1.34% in Twitter-FMNERG) without the prompt-based type-grained queries. This indicates that integrating type-grained information into the query can enhance model performance. Furthermore, adding an additional learnable entity-grained part to the query substantially improves model performance: in the Twitter-GMNER and Twitter-FMNERG datasets, the F1 scores increased by $+2.74\%$ and $+3.18\%$, respectively. The main reason is that these learnable vectors can effectively learn the entity-grained semantics and model intra-entity connections between entity spans and regions, thereby boosting model performance.
The effectiveness of Query-guided Fusion Net. In Table 3, the model without three fusion modules achieves consistent drops. We find the average F1 scores of these three modules decreased by $1.03\%$ and $0.60\%$ on Twitter-GMNER and Twitter-FMNERG, respectively. Experimental results demonstrate that the QFNet can fuse information across different modalities and filter the noise introduced by irrelevant visual regions, which is crucial for improving the performance of two-level relationship modeling.
The effectiveness of Bipartite Matching Loss. In Table 3, excluding the bipartite matching loss from the model leads to an F1 scores decline of $2.18\%$ and $1.76\%$ in Twitter-GMNER and Twitter-FMNERG datasets. This illustrates that the bipartite matching loss contributes to model performance through modeling inter-entity relationships in global matching perspective.
Figure 3: Analysis of candidate visual regions number $k$ for H-Index, TIGER, and our MQSPN.
Discussion and Analysis
Sensitivity Analysis of Irrelevant Visual Regions. The performance of EEG is largely determined by the ground truth coverage of the top-$k$ regions proposed by the class-agnostic RPN. However, a higher value of $k$ implies the introduction of more irrelevant visual regions. To delve into the model’s error tolerance to noisy visual regions, we conduct experiments on H-Index, TIGER, and MQSPN under different values of $k$. As shown in Figure 3, we can observe that the F1 scores of these three methods first increase, and then decrease with the augment of $k$. Compared with H-Index and TIGER, our proposed MQSPN has better and more stable performance on both EEG and GMNER tasks when $k$ becomes larger. This demonstrates MQSPN’s insensitivity to irrelevant visual regions, indicating its excellent error tolerance and robustness of intra-entity connection.
Figure 4: Analysis of the multi-grained queries quantity $u$ on GMNER, MNER and EEG tasks.
Analysis of query quantity. To explore the impact of multi-grained query quantity (i.e., $u$) on GMNER and its 2 subtasks, we report the F1 scores of Twitter-GMNER in Figure 4 by tuning $u$ from $10$ to $90$. We observe that as the number of queries increases from $10$ to $60$, the model’s performance improves by $+1.90\%$, $+1.76\%$, and $+1.45\%$ in the GMNER, MNER, and EEG tasks, respectively. These results indicate that query quantity plays a crucial role in intra-entity connection learning. However, a query quantity exceeding $60$ does not lead to better performance. This suggests that there is an optimal number of queries for MQSPN. In our experiments, we set $u=60$ in the Twitter-GMNER dataset to achieve the optimal results.
Figure 5: (a). Predictions of MQSPN in different ablation setting, which is the same as Table 3. (b). Predictions of TIGER and our MQSPN. The middle and right parts are the visualizations of attention weights in MQSPN. The x-axis denotes the MQS and the y-axis represents sentence tokens or candidate visual regions.
Case Study. We conduct a comprehensive case study in two samples. As shown in Figure 7 (a), when we replace learnable queries with human-designed queries in MQS (w/o LEQ setting), the model cannot differentiate Bradon Saad and Johan Franzen regions, and incorrectly localize Blackhawks and Detroit as Person spans, highlighting deficiencies in intra-entity learning for ambiguous entities. After integrating learnable MQS, the model precisely aligns textual spans with corresponding regions for Person entities. However, using MQS without MSP (w/o BML) leads to confusion in mapping proper inter-entity relationships, resulting in incorrect classifications for Blackhawks and Detroit and more erroneous span-type-region triplets. After combining MQS and MSP, MQSPN effectively differentiates entities and makes accurate predictions, demonstrating the effect of intra-entity and inter-entity modeling.
In Figure 7 (b), TIGER incorrectly identifies the ambiguous entity Jimmy Kimmel within Jimmy Kimmel Live as Person, and associates it with the error region. This results in incorrectly matching the Region-2 with The Weeknd, which originally belongs to Region-0. In contrast, our MQSPN can establish correct intra-entity connections between all textual spans and regions. To further investigate the contribution of MQS, we perform visualizations of the query-text and query-regions attention maps. We find that type-specific queries have the highest attention weights in their target spans or regions, which indicates that type-grained queries of MQS provide effective information. Meanwhile, we observe that different queries are responsible for identifying different spans and regions. For instance, the first PER (person) query mainly focuses on Donald Trump and Region-1, while the second one has the highest attention scores on The Weeknd and Region-0. This demonstrates that these trainable entity-grained queries of MQS adaptively learn distinguishable features for different entities to establish more stable intra-entity connections.
More experiments are attached in the Appendix.
Conclusion
In this paper, we propose a novel method named MQSPN to model appropriate intra-entity and inter-entity relationships for the GMNER task. To achieve this, we propose a learnable Multi-grained Queries Set (MQS) to adaptively learn explicit intra-entity connections. Besides, we reformulate GMNER as a multimodal set prediction (MSP) to model inter-entity relationships from an optimal matching view. Experimental results demonstrate that MQSPN achieves state-of-the-art performance on GMNER and its 2 subtasks across two Twitter benchmarks. In the future, the extension of our MQSPN to the vision-language pre-training model or knowledge-enhanced framework in few-shot or zero-shot scenarios will be one of our directions.
Appendix
Exploration of Performance Upper Bound of MQSPN
For a fair comparison with previous state-of-the-art methods, we only report MQSPN (Bert-base+ViT-B/32+VinVL) in the main paper since it has a similar model scale with TIGER (MMT5-base+VinVL).
Variants of MQSPN Twitter-GMNER Twitter-FMNERG Text Encoder Vision Encoder Region Proposal Network GMNER MNER EEG GMNER MNER EEG Baseline: TIGER (Wang et al. 2023) (MMT5-base + VinVL) 56.63 80.28 61.26 47.20 64.91 61.96 Bert-base ViT-B/32 Faster-RCNN 57.71($\uparrow$1.08) 80.01 61.52 47.86($\uparrow$0.66) 66.83 61.95 Bert-base ViT-B/32 VinVL 58.76($\uparrow$2.13) 80.43 62.40 48.57($\uparrow$1.37) 67.09 62.64 Roberta-large ViT-B/32 VinVL 60.88($\uparrow$4.25) 83.19 63.51 50.39($\uparrow$3.19) 69.64 63.98 Roberta-large ViT-L/14 VinVL 61.12($\uparrow$4.49) 83.24 63.80 51.04($\uparrow$3.82) 69.87 64.35 Fine-tuning Region Proposal Network with Entity Type-related Annotations Bert-base ViT-B/32 Faster-RCNN 56.82 79.71 60.55 47.12 66.31 61.03 Bert-base ViT-B/32 VinVL 57.84 79.94 61.62 47.68 66.67 61.48 Roberta-large ViT-B/32 VinVL 59.90 82.76 62.74 49.53 69.19 63.21 Roberta-large ViT-L/14 VinVL 60.39 82.97 63.28 49.86 69.22 63.67
Table 4: Performance comparison of MQSPN with different foundation models. The first four groups indicate that the Region Proposal Networks (RPN) are derived from their original pretraining checkpoints. VinVL is pre-trained on COCO (Lin et al. 2014), OpenImages (Kuznetsova et al. 2020), Objects365 (Shao et al. 2019), and Visual Genome (Krishna et al. 2017) datasets. Faster-RCNN is pre-trained on PASCAL VOC (Everingham et al. 2010), COCO, and Visual Genome datasets. The last four groups represent that we fine-tune RPNs with entity type-related bounding box annotations.
Actually, MQSPN also achieves higher performance upper bound by employing stronger foundation models as its components. We present the performance of MQSPN with stronger foundation models in GMNER and its two subtasks (i.e. MNER and EEG) across Twitter-GMNER and Twitter-FMNERG datasets, as shown in Table 4. Specifically, we additionally choose xlm-roberta-large [^5] https://hf-mirror.com/FacebookAI/xlm-roberta-large (Roberta-large) for text encoder, clip-vit-large-patch14666https://hf-mirror.com/openai/clip-vit-large-patch14 (ViT-L/14) for vision encoder, and Faster-RCNN [^7] https://github.com/open-mmlab/mmdetection for region proposal network. Different combinations of components are considered to be different variants of MQSPN.
The experimental results from the first four rows illustrate that: (1) Even the weakest MQSPN (Bert-base+ViT-B/32+Faster-RCNN) outperforms the previous state-of-the-art method, TIGER, by $+1.08\%$ and $+0.66\%$ in the Twitter-GMNER and Twitter-FMNERG dataset, respectively, validating our motivation that MQSPN can better enhance the performance of GMNER task by effectively learning intra-entity and inter-entity relationships. (2) Stronger foundation models consistently enhance MQSPN’s performance. The best MQSPN (Roberta-large+ViT-L/14+VinVL) achieves 61.12% and 51.04% F1 scores on the GMNER task for the Twitter-GMNER and Twitter-FMNERG datasets, respectively, which significantly outperforms TIGER by +4.49% and +3.82%. We observe that the text encoder and RPN contribute more significantly to the model’s performance. This is primarily because MQSPN’s strength lies in understanding fine-grained semantic distinction in intra-entity and globally matching inter-entity relations. The text encoder provides robust token-level semantic understanding, while the RPN offers a broader selection of matching regions, supported by region-level feature extraction from the visual backbone.
To further explore the impact of RPNs, we fine-tune our RPNs with entity type-related bounding box annotations. The experimental results are reported in the last four rows of Table 4. We observe that domain-specific fine-tuning of the RPNs cannot result in performance improvement; instead, it leads to a decline. This phenomenon can be mainly attributed to two factors: (1) The bounding box annotations in GMNER are based on fine-grained named entities, leading to the overfitting of RPNs to the training set under this weakly-supervised setting. (2) The entity grounding is inherently an open-world detection problem. Fine-tuning constrains the model to output regions associated with fixed categories, which diminishes the capability of RPNs to generate various candidate regions.
Details of Query-region Prefix Integration
Inspired by prefix tuning (Li and Liang 2021) and its successful applications (Chen et al. 2022a, b) in multimodal fusion, Query-region Prefix Integration is used to reduce the semantic bias between multimodal data. Specifically, we use the query set features as prefix information, inserting them into the candidate region features at keys and values layers of each multi-head attention in the vision transformer (Dosovitskiy et al. 2020). First, the $(l-1)$-th layer candidate region features $H_{V}^{l-1}$ is projected into the query, key, and value vectors :
$q_{V}^{(l)}=H_{V}^{(l-1)}W_{V_{q}}^{(l)},k_{V}^{\left(l\right)}=H_{V}^{\left(l-1\right)}W_{V_{k}}^{\left(l\right)},v_{V}^{\left(l\right)}=H_{V}^{\left(l-1\right)}W_{V_{v}}^{\left(l\right)}$ (13)
$\{\rho_{k}^{l},\rho_{v}^{l}\}=H_{Q}^{l-1}W_{\rho}^{l},$ (14)
$PI^{\left(l\right)}=\mathrm{softmax}(\frac{q_{V}^{\left(l\right)}[\rho_{k}^{\left(l\right)};k_{V}^{\left(l\right)}]^{T}}{\sqrt{h}})[\rho_{v}^{\left(l\right)};v_{V}^{\left(l\right)}]$ (15)
GMNER Datasets